OpenStudy (anonymous):

Medal will be awarded if answer is correct/ accurate and if you explain it cause I really don't get this topic. How many solutions are there to the following system of equations? 4x – 8y = 0 –3x + 8y = –7 A. 0 B. 1 C. infinitely many D. 2

OpenStudy (anonymous):

@ganeshie8 please help just this once

ganeshie8 (ganeshie8):

solve them and see ?

ganeshie8 (ganeshie8):

4x – 8y = 0 –3x + 8y = –7 ----------------------

ganeshie8 (ganeshie8):

add them vertically, wat do u get ? :)

OpenStudy (anonymous):

1x (+ or - idk) and 16y= -7

ganeshie8 (ganeshie8):

y's cancel out, and u wil get value of x

OpenStudy (anonymous):

ok......

ganeshie8 (ganeshie8):

4x – 8y = 0 –3x + 8y = –7 ---------------------- x + 0 = -7 x = -7 you will get some y value for this x, which is obvious. So nothing special about the lines, they will just boringly cut in one point. So one solution

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

One solution x=-7, y=-7/2

OpenStudy (anonymous):

Are you doing linear algebra?

OpenStudy (anonymous):

yea @Moose23

OpenStudy (anonymous):

Thanks again your a lifesaver @ganeshie8

ganeshie8 (ganeshie8):

np :) kind of u lol